Describe the transformations on f(x) that will occur when graphing f(-4x+8).
Answer by robertb(5830)

Case I. f(x) is even.
Then f(-x) = f(x).
==> f(-4x+8) = f(4x-8)
==> f(-4(x-2)) = f(4(x-2))
==> There is horizontal translation 2 units to the right, and then vertical stretching by a factor of 4. There is no reflection across the x-axis.
Case II. f(x) is odd.
Then f(-x) = -f(x).
==> f(-4x+8) = -f(4x-8)
==> f(-4(x-2)) = -f(4(x-2))
==> There is horizontal translation 2 units to the right, and then vertical stretching by a factor of 4. There IS reflection across the x-axis.
Case III. f(x) neither odd nor even is similar to Case II.